# Impact of in-situ nuclear networks and atomic opacities on neutron star merger ejecta dynamics, nucleosynthesis, and kilonovae

### v1

@@ 4 Conclusions
In this paper, we carried out a systematic analysis of the impact of fully coupling nuclear calculations to the hydrodynamic evolution of ejecta from BNSM on their dynamics, nucleosynthesis, and kilonova light curves. We evolved ejecta profiles extracted from ab-initio NR simulations up to $t\sim 30$ days, and explored different nuclear burning, thermalization, and opacity models. Our most advanced setup couples the information provided by the in-situ NN to composition-dependent, nuclear-informed thermalization calculations and atomic-physics based, frequency-dependent opacities.
@@
The in-situ NN qualitatively impacts the early-time ($t\sim 1$ hour) kilonova light curves by altering the location and temperature of the photosphere. As a result, the emission is dimmer, redder, and peaks later compared to models using analytical fits for the nuclear power. These simpler prescriptions perform relatively better at later times in approximating the light curves. However, employing an in-situ NN also enables the use of composition-dependent thermalization and opacity models, whose significant qualitative effects on the EM emission are summarized in the remainder of this section. The global corrections introduced by the online nuclear calculations to the density evolution of the outflow have negligible dynamical effects.
@@
The overall improvements enabled by the introduction of an in-situ NN can be summarized as follows. Resolving the detailed hydrodynamical evolution of the outflow during the first several hundreds milliseconds is essential for obtaining reliable nucleosynthesis predictions. Calibrated analytical fits for the nuclear power, together with constant prescriptions for the thermalization efficiency and optical opacity, are sufficient to capture the bulk dynamics of the outflow. However, a more accurate thermalization scheme is required to correctly reproduce its temperature evolution. Moreover, we find that both the composition-dependent thermalization scheme and the multi-group opacity treatment are necessary to accurately model the kilonova emission.
@@
In this paper, we used a Lagrangian approach and we did not investigate the role of composition mixing. Some degree of mixing is physically expected, and its relevance is set by the Damköler number $D_{\alpha}$, expressing the ratio between the mixing and the (nuclear) reaction timescales (Dimotakis, 2005). For $D_{\alpha}\sim 1$, composition mixing and nuclear reactions cannot be decouple and might affect each other. Lagrangian codes usually assumes $D_{\alpha}\gg 1$ (mixing is not relevant), while Eulerian calculations typically overestimate the impact of mixing due to numerical dissipation. In the future, we plan to investigate how a flow of isotopes between neighboring fluid elements can affect the nucleosynthesis calculation and the resulting heating rates exploring different regimes for the mixing length.
@@ Appendix A Radiative transport equations
The transport equation for the specific intensity $I_{\nu}(\@vec{r},\@vec{\Omega},t)$ of radiation has the form
$\frac{1}{c}\frac{\partial I_{\nu}}{\partial t}+\@vec{n}\cdot\nabla I_{\nu}+\sigma_{\nu}I_{\nu}=\frac{c}{4\pi}(\sigma_{\nu}^{a}\epsilon_{\nu}^{P}-\sigma_{\nu}^{s}\epsilon_{\nu}^{r})\,$, (24)
where $c$ is the speed of light, n\@vec{n} is the direction of propagation, and $\sigma_{\nu}^{a}(\@vec{r},t)=\kappa_{\nu}\rho$, with $\rho$ density and $\kappa_{\nu}$ opacity, and $\sigma_{\nu}^{s}(\@vec{r},t)$ are the absorption and scattering coefficients, with $\sigma_{\nu}=\sigma_{\nu}^{a}+\sigma_{\nu}^{s}$. The emissivity is $\epsilon_{\nu}^{r}(\@vec{r},t)\equiv c^{-1}\int_{4\pi}d\Omega I_{\nu}$, with $\Omega$ solid angle, and $\epsilon_{\nu}^{P}=4\pi B_{\nu}/c$ is the blackbody energy density. Integrating Eq. (24) over $d\Omega$ gives
@@
Local thermodynamical equilibrium (LTE) is a sufficient (since it implies IDA) but not necessary condition for this approximation, for which only isotropic radiation on distances smaller than the mean free path $(\sigma_{\nu}^{a})^{-1}$ of the radiation is required. If LTE, then $\epsilon_{\nu}^{P}=\epsilon_{\nu}^{r}$ and thus
@@
$\frac{1}{c}\frac{\partial(\epsilon_{\nu}^{r}\psi_{\nu})}{\partial t}+\@vec{n}\cdot\nabla(\epsilon_{\nu}^{r}\psi_{\nu})+\sigma_{\nu}\epsilon_{\nu}^{r}\psi_{\nu}=\frac{1}{4\pi}(\sigma_{\nu}^{a}\epsilon_{\nu}^{P}-\sigma_{\nu}^{s}\epsilon_{\nu}^{r})\,.$ (30)
Its angular integral still gives Eq. (25), as $\psi_{\nu}$ is normalized as $\int_{4\pi}d\Omega\psi_{\nu}=1$. Using only the assumption of a slowing varying normalized intensity, $\partial_{t}\psi_{\nu}+c\@vec{n}\cdot\nabla\psi_{\nu}=0$, one can find
@@
$\@vec{R}_{\nu}=-\frac{\nabla\epsilon_{\nu}^{r}}{\sigma_{\nu}\omega\epsilon_{\nu}^{r}}\,$, (32)
@@
In IDA, $R_{\nu}\to 0$, $\psi_{\nu}\to(4\pi)^{-1}$ and $\lambda_{\nu}\to 1$. Under the LTE assumption, $\epsilon_{\nu}^{r}=\epsilon_{\nu}^{P}$, $\omega=1$, and
@@
$\@vec{F}_{\nu}=-\frac{\lambda_{\nu}c}{\sigma_{\nu}}\nabla\epsilon_{\nu}^{P}=-\frac{4\pi\lambda_{\nu}}{3\kappa_{\nu}\rho}\frac{\partial B_{\nu}}{\partial T}\nabla T\,$, (36)
where $\sigma_{\nu}=\kappa_{\nu}\rho$, with $\kappa_{\nu}$ total opacity. In spherical symmetry (a single ray-by-ray section is an effective 1D problem), $\nabla T=dT/dr=4\pi r^{2}\rho\,dT/dm$, and thus the flux and flux limiter can be written as in Eq. (3)-(5).
@@ Appendix B Network Initialization
To investigate the systematic uncertainties on the nucleosynthesis arising from the NSE-based initialization procedures, we perform the same simulation three times, initializing the nuclear composition either with the backtracking method at $T_{\text{NSE}}=6,8$ GK (BK6 and BK8), or directly at the temperature recorded at extraction (CNSE).
In Fig. 18, we show the composition for the BLh ejecta profile at the beginning of the simulation, during the first phase of $r$-process nucleosynthesis, and once the production of the heavy nuclei is complete. The cold-NSE initialization misses the early reactions after NSE drop-out (physically happening for many fluid elements during NR simulation times, $t<0$). Moreover, the low temperatures at extraction cause the Boltzmann term in the NSE equation to be overestimated (see Eq. (8)). Together, these two effects suppress or inflate the initial abundances of the isotopes around $A\sim 65$ and $A\sim 80$, respectively, and fail to capture the very early production of elements above the first peak. The BK6 and BK8 initializations are considerably more similar to each other, with differences from the lower temperature qualitatively similar to those discussed for the cold-NSE approach. After initialization and during the $r$-process nucleosynthesis ($1\penalty 10000\ \mathrm{ms}\lesssim t\lesssim 1$ s), the rates of the dominant nuclear reactions are set by the local composition. These rates evolve in a way that counteracts the initial errors, progressively reducing the systematic uncertainty over time. As a result, the system gradually loses memory of the initialization method, with relative discrepancies in the late-time yields falling below unity for most mass numbers. As a notable exception, a group of isotopes slightly lighter than the iron group is overproduced by the more approximated initialization methods.
Figure 18: Top panels: Mass-weighted composition of the BLh ejecta profile after initialization (left column), at $t=0.1$ s (pre neutron freeze-out, central column), and $t=5\times 10^{4}$ s (right column). Blue, orange, and green lines show results respectively from the backtrack ($T_{\text{NSE}}=8$ GK and $T_{\text{NSE}}=6$ GK) and the cold-NSE initialization procedures. The black dots represent the Solar r-abundances (Prantzos et al. 2020). The abundances are scaled to match a unitary cumulative abundance for the elements of the first $r$-process peak, with $72\leq A\leq 94$. The vertical dashed lines highlight the position of the elements $A=56$. The blue, orange, and purple shaded regions remark the position of the first, second and third $r$-process peaks, respectively. Bottom panel: relative differences against the BK8 run. The upwards [downwards] triangles indicate discrepancies greater than two orders of magnitude [smaller than $5\times 10^{-4}$]. The horizontal dashed lines mark the factors $0.01$, 1, and 10 discrepancies.
Single isotopes can experience more significant discrepancies. The cold-NSE method tends to leave less free neutrons after freeze-out, due to a stronger initial presence of seed nuclei. In our case, this introduces only minor corrections to the light curves, and only for $t\lesssim 0.5$ days.
@@
The backtracking at the higher temperature $T_{\text{NSE}}=8$ GK has in general to be preferred over the alternatives, as it approximately takes into account the otherwise overlooked early out-of-NSE processes. However, backtracking initialization methods assume that the pre-evolution is short enough ($t\ll 1$ s) to have an approximately negligible effect on the electron fraction and entropy. We check these hypotheses in Fig. 19. For the high initialization temperature, around $90\%$ of the ejected mass shows relative discrepancies respectively on the electron fraction and entropy below $10\%$ or $3\%$ after the pre-evolution. For around $75\%$ of the mass, this evolution lasts less than $30$ ms. The backtracking assumptions are better satisfied by the low-temperature case. For most of the ejecta, the backtracking is safe and closes before the ignition of the core reactions of the $r$-process nucleosynthesis. Few percent of the fluid elements register pre-evolution times of $30$ ms or more, and a change in the $Y_{e}$ higher than $10\%$. Such a discrepancy can be crucial for accessing the strong $r$-process nucleosynthesis for $Y_{e}\sim 0.22$ and to distinguish between neutron and proton rich material at $Y_{e}\sim 0.5$. Long pre-evolution times also mean missing the NN-hydrodynamics coupling in the first phases of the $r$-process nucleosynthesis. Figure 18 provides an estimate of the systematic uncertainties propagated from our initialization methods. In the future, we plan to reduce these uncertainties by initializing the fluid element compositions using tracer particles extracted from the original NR data and pre-processed up to the initial time of the hydrodynamic simulation.

### v2

@@ 4 Conclusions
In this paper, we carried out a systematic analysis of the impact of fully coupling nuclear calculations to the hydrodynamic evolution of ejecta from BNSMs on their dynamics, nucleosynthesis, and kilonova light curves. We evolved ejecta profiles extracted from ab-initio NR simulations up to $t\sim 30$ days, and explored different nuclear burning, thermalization, and opacity models. Our most advanced setup couples the information provided by the in-situ NN to composition-dependent, nuclear-informed thermalization calculations and atomic-physics based, frequency-dependent opacities.
@@
The in-situ NN qualitatively impacts the early-time ($t\sim 1$ hour) kilonova light curves by altering the location and temperature of the photosphere. As a result, the emission is dimmer, redder, and peaks later compared to models using analytical fits for the nuclear power. These simpler prescriptions perform relatively better in approximating the light curves at later times. However, employing an in-situ NN also enables the use of composition-dependent thermalization and opacity models, whose significant qualitative effects on the EM emission are summarized in the remainder of this section. The global corrections introduced by the online nuclear calculations to the density evolution of the outflow have negligible dynamical effects.
@@
The overall improvements enabled by the introduction of an in-situ NN can be summarized as follows. Resolving the detailed hydrodynamical evolution of the outflow during the first several hundred milliseconds is essential for obtaining reliable nucleosynthesis predictions. Calibrated analytical fits for the nuclear power, together with constant prescriptions for the thermalization efficiency and optical opacity, are sufficient to capture the bulk dynamics of the outflow. However, a more accurate thermalization scheme is required to correctly reproduce its temperature evolution. Moreover, we find that both the composition-dependent thermalization scheme and the multigroup opacity treatment are necessary to accurately model the kilonova emission.
@@
In this paper, we used a Lagrangian approach and we did not investigate the role of composition mixing. Some degree of mixing is physically expected, and its relevance is set by the Damköler number $D_{\alpha}$, expressing the ratio between the mixing and the (nuclear) reaction timescales (Dimotakis, 2005). For $D_{\alpha}\sim 1$, composition mixing and nuclear reactions cannot be decoupled and might affect each other. Lagrangian codes usually assume $D_{\alpha}\gg 1$ (mixing is not relevant), while Eulerian calculations typically overestimate the impact of mixing due to numerical dissipation. In Zhai et al. (2026), the effect of radial mixing is found to negligibly affect the global nucleosynthesis and light curves. In the future, we plan to deepen this analysis using the more advanced opacity models introduced in this paper and to extend it to simulations in higher dimensions.
@@ Appendix A Radiative transport equations
For matter in LTE (but not necessarily in equilibrium with radiation), the transport equation for the specific intensity $I_{\nu}(\@vec{r},\@vec{\Omega},t)$ of radiation has the form (e.g. Levermore & Pomraning 1981)
$\frac{1}{c}\frac{\partial I_{\nu}}{\partial t}+\@vec{n}\cdot\nabla I_{\nu}+\sigma_{\nu}I_{\nu}=\frac{c}{4\pi}(\sigma_{\nu}^{a}\epsilon_{\nu}^{P}+\sigma_{\nu}^{s}\epsilon_{\nu}^{r})\,$, (24)
where $c$ is the speed of light, n\@vec{n} is the direction of propagation, $\sigma_{\nu}^{a,s}(\@vec{r},t)$ are the absorption and scattering coefficients, with $\sigma_{\nu}=\sigma_{\nu}^{a}+\sigma_{\nu}^{s}$ and $\sigma_{\nu}=\kappa_{\nu}\rho$, $\rho$ is the density, and $\kappa_{\nu}$ is the opacity. The specific energy density of the radiation field is $\epsilon_{\nu}^{r}(\@vec{r},t)\equiv c^{-1}\int_{4\pi}d\Omega I_{\nu}$, with $\Omega$ solid angle, and $\epsilon_{\nu}^{P}=4\pi B_{\nu}/c$ is the blackbody energy density. Integrating Eq. (24) over $d\Omega$ gives
@@
Radiation-matter equilibrium is a sufficient (since it implies IDA) but not necessary condition for this approximation, for which only isotropic radiation on distances smaller than the mean free path $(\sigma_{\nu}^{a})^{-1}$ of the radiation is required. If radiation-matter equilibrium is assumed (single-temperature diffusion approximation), then $\epsilon_{\nu}^{P}=\epsilon_{\nu}^{r}$ and thus
@@
$\frac{1}{c}\frac{\partial(\epsilon_{\nu}^{r}\psi_{\nu})}{\partial t}+\@vec{n}\cdot\nabla(\epsilon_{\nu}^{r}\psi_{\nu})+\sigma_{\nu}\epsilon_{\nu}^{r}\psi_{\nu}=\frac{1}{4\pi}(\sigma_{\nu}^{a}\epsilon_{\nu}^{P}+\sigma_{\nu}^{s}\epsilon_{\nu}^{r})\,.$ (30)
Its angular integral still gives Eq. (25), as $\psi_{\nu}$ is normalized as $\int_{4\pi}d\Omega\psi_{\nu}=1$. Using only the assumption of a slowly varying normalized intensity, $\partial_{t}\psi_{\nu}+c\@vec{n}\cdot\nabla\psi_{\nu}=0$, one can find
@@
$\@vec{R}_{\nu}=-\frac{\nabla\epsilon_{\nu}^{r}}{\sigma_{\nu}\omega_{\nu}\epsilon_{\nu}^{r}}\,$, (32)
@@
In IDA, $R_{\nu}\to 0$, $\psi_{\nu}\to(4\pi)^{-1}$ and $\lambda_{\nu}\to 1$. In the single-temperature diffusion approximation, $\epsilon_{\nu}^{r}=\epsilon_{\nu}^{P}$, $\omega=1$, and
@@
$\@vec{F}_{\nu}=-\frac{\lambda_{\nu}c}{3\sigma_{\nu}}\nabla\epsilon_{\nu}^{P}=-\frac{4\pi\lambda_{\nu}}{3\kappa_{\nu}\rho}\frac{\partial B_{\nu}}{\partial T}\nabla T\,.$ (36)
In spherical symmetry (a single ray-by-ray section is an effective 1D problem), $\nabla T=dT/dr=4\pi r^{2}\rho\,dT/dm$, and thus the flux and flux limiter can be written as in Eq. (3)-(5).
@@ Appendix B Network initialization
Figure 19: Top panels: Mass-weighted composition of the BLh ejecta profile after initialization (left column), at $t=0.1$ s (pre neutron freeze-out, central column), and $t=5\times 10^{4}$ s (right column). Blue, orange, and green lines show results respectively from the backtrack ($T_{\text{NSE}}=8$ GK and $T_{\text{NSE}}=6$ GK) and the cold-NSE initialization procedures. The black dots represent the Solar r-abundances (Prantzos et al. 2020). The abundances are scaled to match a unitary cumulative abundance for the elements of the first $r$-process peak, with $72\leq A\leq 94$. The vertical dashed lines highlight the position of the elements $A=56$. The blue, orange, and purple shaded regions remark the position of the first, second and third $r$-process peaks, respectively. Bottom panel: relative differences against the BK8 run. The upward [downward] triangles indicate discrepancies greater than two orders of magnitude [smaller than $5\times 10^{-4}$]. The horizontal dashed lines mark the factors $0.01$, 1, and 10 discrepancies.
@@
To investigate the systematic uncertainties on the nucleosynthesis arising from the NSE-based initialization procedures, we perform the same simulation three times, initializing the nuclear composition either with the backtracking method at $T_{\text{NSE}}=6,8$ GK (BK6 and GK simulations), or directly at the temperature recorded at extraction (CNSE).
In Fig. 19, we show the composition for the BLh ejecta profile at the beginning of the simulation, during the first phase of $r$-process nucleosynthesis, and once the production of the heavy nuclei is complete. The cold-NSE initialization misses the early reactions after NSE drop-out (physically happening for many fluid elements during NR simulation times, $t<0$). Moreover, the low temperatures at extraction cause the Boltzmann term in the NSE equation to be overestimated (see Eq. (8)). Together, these two effects suppress or inflate the initial abundances of the isotopes around $A\sim 65$ and $A\sim 80$, respectively, and fail to capture the very early production of elements above the first peak. The BK6 and BK8 initializations are considerably more similar to each other, with differences from the lower temperature qualitatively similar to those discussed for the cold-NSE approach. After initialization and during the $r$-process nucleosynthesis ($1\penalty 10000\ \mathrm{ms}\lesssim t\lesssim 1$ s), the rates of the dominant nuclear reactions are set by the local composition. These rates evolve in a way that counteracts the initial errors, progressively reducing the systematic uncertainty over time. As a result, the system gradually loses memory of the initialization method, with relative discrepancies in the late-time yields falling below unity for most mass numbers. This effect is reflected in the final lanthanides mass fractions reported in Table 4, which agree within a few percent across the GK, BK6, and CNSE simulations. As a notable exception, a group of isotopes slightly lighter than the iron group is overproduced by the more approximated initialization methods. Single isotopes can experience more significant discrepancies. The cold-NSE method tends to leave less free neutrons after freeze-out, due to a stronger initial presence of seed nuclei. In our case, this introduces only minor corrections to the light curves, and only for $t\lesssim 0.5$ days.
The backtracking at the higher temperature $T_{\text{NSE}}=8$ GK has in general to be preferred over the alternatives, as it approximately takes into account the otherwise overlooked early out-of-NSE processes. However, backtracking initialization methods assume that the pre-evolution is short enough ($t\ll 1$ s) to have an approximately negligible effect on the electron fraction and entropy. We check these hypotheses in Fig. 20. For the high initialization temperature, around $90\%$ of the ejected mass shows relative discrepancies respectively on the electron fraction and entropy below $10\%$ or $3\%$ after the pre-evolution. For around $75\%$ of the mass, this evolution lasts less than $30$ ms. The backtracking assumptions are better satisfied by the low-temperature case. For most of the ejecta, the backtracking is safe and closes before the ignition of the core reactions of the $r$-process nucleosynthesis. Few percent of the fluid elements register pre-evolution times of $30$ ms or more, and a change in the $Y_{e}$ higher than $10\%$. Such a discrepancy can be crucial for accessing the strong $r$-process nucleosynthesis for $Y_{e}\sim 0.22$ and to distinguish between neutron and proton rich material at $Y_{e}\sim 0.5$. Long pre-evolution times also mean missing the NN-hydrodynamics coupling in the first phases of the $r$-process nucleosynthesis. Figure 19 provides an estimate of the systematic uncertainties propagated from our initialization methods. In the future, we plan to reduce these uncertainties by initializing the fluid element compositions using tracer particles extracted from the original NR data and preprocessed up to the initial time of the hydrodynamic simulation.